# ReQISC: A Reconfigurable Quantum Computer Microarchitecture and Compiler Co-Design

### v1

@@ A.4. Canonicalization & Reparameterization for genAshN-EA
$\small H_{EA}:=a(\sigma_{X}\otimes\sigma_{X}+\sigma_{Y}\otimes\sigma_{Y}+\sigma_{Z}\otimes\sigma_{Z})-\eta\cdot\sigma_{Y}\otimes\sigma_{Y}-\sigma_{Z}\otimes\sigma_{Z}+\Omega(\sigma_{X}\otimes I+I\otimes\sigma_{X})+\delta(\sigma_{Z}\otimes I+I\otimes\sigma_{Z}).$
@@
$P(\lambda)$ $:=\frac{\det(H_{EA}-\lambda I)}{\lambda+3a-1-\eta}.$
@@ A.5. Correspondence between $(\alpha,\beta)$, $T$ and $(x,y,z)$
With the above reparameterization, we know explicitly the eigenspectrum of $H_{EA}$. We can therefore write down the following function of the trace of
$V(a,a-\eta,a-1;\tau;\sqrt{(1-\alpha)\beta(1-\eta+\alpha+\beta)},0,\sqrt{\alpha(1+\beta)(\beta+\alpha-\eta)})$:
@@
$\small(\Omega,\delta)\in\mathbb{R}^{2}_{\geq 0}\xrightarrow{\phi_{1}}(\alpha,\beta)\in Q_{\eta}\xrightarrow{\phi_{2}}T\in S\xrightarrow{\phi_{3}}(x,y,z)\in EA_{+}.$
@@
$\small\begin{cases}x_{i}^{-1}=\bar{x}_{i},i=1,2,3,\\ x_{1}x_{2}x_{3}=-\exp\{i\tau(1+\eta-3a)\},\\ x_{1}+x_{2}+x_{3}=e^{-ia\tau}T.\end{cases}$
@@
$=$ $x^{3}-(x_{1}+x_{2}+x_{3})x^{2}+(x_{1}x_{2}+x_{1}x_{3}+x_{2}x_{3})x-x_{1}x_{2}x_{3}$
$=$ $x^{3}-e^{-ia\tau}Tx^{2}-\exp\{i\tau(1+\eta-3a)\}((x_{3}^{-1}+x_{2}^{-1}+x_{1}^{-1})x-1)$
$=$ $x^{3}-e^{-ia\tau}Tx^{2}-\exp\{i\tau(1+\eta-3a)\}(\overline{(x_{1}+x_{2}+x_{3})}x-1)$
$=$ $x^{3}-e^{-ia\tau}Tx^{2}-\exp\{i\tau(1+\eta-3a)\}(e^{ia\tau}\bar{T}x-1),$
@@ A.6.1. Continuity of $\phi_{2}$
$TG$ $=\frac{\gamma\left(\alpha\gamma(\eta-1)+\alpha-\gamma\eta^{2}+\gamma+\eta-1\right)e^{-i\tau\left(-2\alpha-\frac{2}{\gamma}+\eta+1\right)}}{(\gamma(\alpha-\eta-1)+2)(\gamma(2\alpha-\eta-1)+1)}+\frac{\gamma(\alpha\gamma(\alpha-\eta-1)+\alpha+\eta-1)e^{\frac{i\tau(\gamma\eta+\gamma-2)}{\gamma}}}{(\alpha\gamma-1)(\gamma(\alpha-\eta-1)+2)}$
(6) $+\frac{(\gamma(\alpha(\gamma(-\eta)+\gamma-1)+\eta+1)-1)e^{i\tau(-2\alpha+\eta+1)}}{(\alpha\gamma-1)(\gamma(2\alpha-\eta-1)+1)}$
@@
$\gamma(\alpha-\eta-1)+2=0$
$\gamma(2\alpha-\eta-1)+1=0$
@@
$\alpha+2/\gamma-1=\eta$
$2\alpha+1/\gamma-1=\eta$
@@
$\alpha\to 0,\gamma\to 1/(1+\eta)$
$\alpha\to 1,\gamma\to 1$
@@
$TG|_{\alpha=\epsilon,\gamma=1/(1+\eta)+h}$ $=e^{i\tau(1+\eta)}\frac{2(\eta+1)(\eta(h-1)+h)}{(\eta+(1+\eta)h(-\eta+\epsilon-1)+\epsilon+1)(-\eta+\epsilon(\eta h+h+1)-1)}$
$+e^{-i\tau(1+\eta)}O(1)+O(\epsilon+h).$
@@
$TG|_{\alpha=1-\epsilon,\gamma=1-\epsilon^{\prime}}$ $=-e^{-i\tau(-4+\eta+1)}\frac{\left(\epsilon^{\prime}-1\right)\left((\eta-1)(\eta+\epsilon)\epsilon^{\prime}-\eta(\eta+\epsilon-2)\right)}{\left(-\eta+(\eta+\epsilon)\epsilon^{\prime}-\epsilon+2\right)\left(-\eta+(\eta+2\epsilon-1)\epsilon^{\prime}-2\epsilon+2\right)}$
$+e^{-i\tau(1-\eta)}\frac{2\left(-\eta+(\eta+\epsilon-1)\epsilon^{\prime}-\epsilon+2\right)\left(-\eta+(\eta+\epsilon)\epsilon^{\prime}-\epsilon+1\right)}{\left(-\eta+(\eta+\epsilon)\epsilon^{\prime}-\epsilon+2\right)\left(-\eta+(\eta+2\epsilon-1)\epsilon^{\prime}-2\epsilon+2\right)}+O(\epsilon+h).$
@@ A.6.2. Surjectivity of $\phi_{2}$
$TG|_{\alpha=0}$ $=\frac{e^{\frac{i\tau(\gamma\eta+\gamma-2)}{\gamma}}\left(\gamma(\eta-1)\left(-e^{-\frac{2i\tau(\gamma\eta+\gamma-2)}{\gamma}}\right)+(\gamma\eta+\gamma-2)e^{\frac{2i\tau}{\gamma}}+\gamma(\eta-1)\right)}{\gamma\eta+\gamma-2},$
$TG|_{\alpha=1}$ $=\frac{e^{-\frac{i\tau(\gamma(\eta-1)-2)}{\gamma}}\left(\gamma\eta e^{\frac{2i\tau(\gamma\eta-2)}{\gamma}}+(\gamma\eta-2)e^{\frac{2i\tau(\gamma(\eta-1)-1)}{\gamma}}-\gamma\eta\right)}{\gamma\eta-2},$
$TG|_{\gamma=1}$ $=-\frac{e^{i(\eta-1)\tau}\left(2\alpha+\eta\left(e^{2i\tau(\alpha-\eta+1)}-e^{-2i(\alpha-1)\tau}-1\right)\right)}{\eta-2\alpha},$
@@
$TG|_{\alpha=-1/\gamma+1+\eta}$ $=\frac{e^{\frac{i\tau(\gamma\eta+\gamma-2)}{\gamma}}\left(\gamma(\eta-1)\left(-e^{-\frac{2i\tau(\gamma\eta+\gamma-2)}{\gamma}}\right)+(\gamma\eta+\gamma-2)e^{\frac{2i\tau}{\gamma}}+\gamma(\eta-1)\right)}{\gamma\eta+\gamma-2}.$
We consider the contour $(\eta,1)-(0,1/(1+\eta))-(0,0)-(1,0)-(1,1)-(\eta,1)$ around $U_{\eta}$. We compute
$TG|_{(\alpha,\gamma)=(\eta,1)}$ $=e^{i\tau(\eta-1)}+e^{i\tau(\eta+1)}-e^{i\tau(1-\eta)}=A$
$B:=TG|_{(\alpha,\gamma)=(0,1/(1+\eta))}$ $=e^{-i\tau(1+\eta)}\frac{1-\eta}{1+\eta}+e^{i\tau(1+\eta)}\frac{2\eta}{1+\eta}$
$TG|_{(\alpha,\gamma)=(0,0)}$ $=e^{i\tau(\eta+1)}=C$
$TG|_{(\alpha,\gamma)=(1,0)}$ $=e^{i\tau(\eta-1)}=D$
$E:=TG|_{(\alpha,\gamma)=(1,1)}$ $=e^{i\tau(\eta-1)}2\frac{1-\eta}{2-\eta}+e^{-i\tau(\eta-3)}\frac{\eta}{2-\eta}$
@@
$TG|_{\alpha=0}$ $=-e^{-i\tau(\eta+1-2/\gamma)}\frac{\gamma(\eta-1)}{\gamma(\eta+1)-2}+e^{i\tau(\eta+1)}+e^{i\tau(\eta+1-2/\gamma)}\frac{\gamma(\eta-1)}{\gamma(\eta+1)-2}$
$=e^{i\tau(\eta+1)}+2(\eta-1)\frac{\sin[\tau(\eta+1-2/\gamma)]}{\eta+1-2/\gamma}\times i,$
@@
$TG|_{\gamma=1}=e^{i\tau(\eta-1)}+\frac{2\eta}{\eta-2\alpha}\sin[\tau(\eta-2\alpha)]\times ie^{i\tau}.$
@@
$TG|_{\alpha=1}=e^{i\tau(\eta-1)}+\frac{2\gamma\eta}{\gamma\eta-2}\sin[\tau(\eta-2/\gamma)]\times ie^{i\tau},$
@@ A.6.3. Surjectivity of $\phi_{3}$
$\phi_{3}^{-1}:P$ $\mapsto e^{i\tau(\eta+1)}-e^{i\tau(1-\eta)}+e^{i\tau(\eta-1)}=A$
$Q$ $\mapsto e^{i\tau(\eta+1)}-1+1=e^{i\tau(\eta+1)}=C$
$R$ $\mapsto e^{i\tau(\eta+1)/3}$

### v2

@@ A.4. Canonicalization & Reparameterization for genAshN-EA
$H_{EA}$ $:=a(\sigma_{X}\otimes\sigma_{X}+\sigma_{Y}\otimes\sigma_{Y}+\sigma_{Z}\otimes\sigma_{Z})$
$\quad-\eta\cdot\sigma_{Y}\otimes\sigma_{Y}-\sigma_{Z}\otimes\sigma_{Z}$
$\quad+\Omega(\sigma_{X}\otimes I+I\otimes\sigma_{X})$
(40) $\quad+\delta(\sigma_{Z}\otimes I+I\otimes\sigma_{Z}).$
@@
(41) $P(\lambda):=\frac{\det(H_{EA}-\lambda I)}{\lambda+3a-1-\eta}.$
@@ A.5. Correspondence between $(\alpha,\beta)$, $T$ and $(x,y,z)$
Given the aforementioned reparameterization, the eigenspectrum of $H_{EA}$ can be determined explicitly. This allows us to derive the trace of $V(a,a-\eta,a-1;\tau;\Omega,0,\delta)$ as a function of the parameters, where $\Omega=\sqrt{(1-\alpha)\beta(1-\eta+\alpha+\beta)}$ and $\delta=\sqrt{\alpha(1+\beta)(\beta+\alpha-\eta)}$:
@@
$(\Omega,\delta)\in\mathbb{R}^{2}_{\geq 0}\xrightarrow{\phi_{1}}(\alpha,\beta)\in Q_{\eta}\xrightarrow{\phi_{2}}T\in S\xrightarrow{\phi_{3}}(x,y,z)\in EA_{+}.$
@@
(43) $\begin{cases}x_{i}^{-1}=\bar{x}_{i},i=1,2,3,\\ x_{1}x_{2}x_{3}=-\exp\{i\tau(1+\eta-3a)\},\\ x_{1}+x_{2}+x_{3}=e^{-ia\tau}T.\end{cases}$
@@
$=x^{3}-(x_{1}+x_{2}+x_{3})x^{2}$
$\quad+(x_{1}x_{2}+x_{1}x_{3}+x_{2}x_{3})x-x_{1}x_{2}x_{3}$
$=x^{3}-e^{-ia\tau}Tx^{2}-\exp\{i\tau(1+\eta-3a)\}$
$\quad\times((x_{3}^{-1}+x_{2}^{-1}+x_{1}^{-1})x-1)$
$=x^{3}-e^{-ia\tau}Tx^{2}-\exp\{i\tau(1+\eta-3a)\}$
(44) $\quad\times(e^{ia\tau}\bar{T}x-1),$
@@ A.6.1. Continuity of $\phi_{2}$
$TG$ $=\frac{\gamma(\alpha\gamma(\eta-1)+\alpha-\gamma\eta^{2}+\gamma+\eta-1)}{(\gamma(\alpha-\eta-1)+2)(\gamma(2\alpha-\eta-1)+1)}$
$\qquad\times e^{-i\tau(-2\alpha-2/\gamma+\eta+1)}$
$\,+\frac{\gamma(\alpha\gamma(\alpha-\eta-1)+\alpha+\eta-1)}{(\alpha\gamma-1)(\gamma(\alpha-\eta-1)+2)}$
$\qquad\times e^{i\tau(\gamma\eta+\gamma-2)/\gamma}$
$\,+\frac{(\gamma(\alpha(\gamma(-\eta)+\gamma-1)+\eta+1)-1)}{(\alpha\gamma-1)(\gamma(2\alpha-\eta-1)+1)}$
(46) $\qquad\times e^{i\tau(-2\alpha+\eta+1)}$
@@
$\gamma(\alpha-\eta-1)+2$ $=0,$
(49) $\gamma(2\alpha-\eta-1)+1$ $=0,$
@@
$\alpha+2/\gamma-1$ $=\eta,$
(50) $2\alpha+1/\gamma-1$ $=\eta,$
@@
(51) $\alpha\to 0,\,\gamma\to 1/(1+\eta)\quad\text{and}\quad\alpha\to 1,\,\gamma\to 1$
@@
$TG|_{\alpha=\epsilon,\gamma=1/(1+\eta)+h}$
$=e^{i\tau(1+\eta)}\frac{2(\eta+1)(\eta(h-1)+h)}{(\eta+(1+\eta)h(-\eta+\epsilon-1)+\epsilon+1)}$
$\quad\times\frac{1}{(-\eta+\epsilon(\eta h+h+1)-1)}$
(52) $\quad+e^{-i\tau(1+\eta)}O(1)+O(\epsilon+h).$
@@
$TG|_{\alpha=1-\epsilon,\gamma=1-\epsilon^{\prime}}$
$=-e^{-i\tau(-4+\eta+1)}\frac{(\epsilon^{\prime}-1)((\eta-1)(\eta+\epsilon)\epsilon^{\prime}-\eta(\eta+\epsilon-2))}{(-\eta+(\eta+\epsilon)\epsilon^{\prime}-\epsilon+2)}$
$\quad\times\frac{1}{(-\eta+(\eta+2\epsilon-1)\epsilon^{\prime}-2\epsilon+2)}$
$\quad+e^{-i\tau(1-\eta)}\frac{2(-\eta+(\eta+\epsilon-1)\epsilon^{\prime}-\epsilon+2)}{(-\eta+(\eta+\epsilon)\epsilon^{\prime}-\epsilon+2)}$
(53) $\quad\times\frac{(-\eta+(\eta+\epsilon)\epsilon^{\prime}-\epsilon+1)}{(-\eta+(\eta+2\epsilon-1)\epsilon^{\prime}-2\epsilon+2)}+O(\epsilon+h).$
@@ A.6.2. Surjectivity of $\phi_{2}$
$TG|_{\alpha=0}$ $=\frac{e^{A_{1}}\bigl[\gamma(\eta-1)(1-e^{-2A_{1}})+(\gamma\eta+\gamma-2)e^{A_{2}}\bigr]}{\gamma\eta+\gamma-2},$
$TG|_{\alpha=1}$ $=\frac{e^{-A_{3}}\bigl[\gamma\eta e^{A_{4}}+(\gamma\eta-2)e^{A_{5}}-\gamma\eta\bigr]}{\gamma\eta-2},$
$TG|_{\gamma=1}$ $=-\frac{e^{i(\eta-1)\tau}\bigl[2\alpha+\eta\bigl(e^{2i\tau(\alpha-\eta+1)}-e^{-2i(\alpha-1)\tau}-1\bigr)\bigr]}{\eta-2\alpha},$
@@
$TG|_{\alpha=-\frac{1}{\gamma}+1+\eta}$ $=\frac{e^{A_{1}}\bigl[\gamma(\eta-1)(1-e^{-2A_{1}})+(\gamma\eta+\gamma-2)e^{A_{2}}\bigr]}{\gamma\eta+\gamma-2},$
where $A_{1}=\frac{i\tau(\gamma\eta+\gamma-2)}{\gamma}$, $A_{2}=\frac{2i\tau}{\gamma}$, $A_{3}=\frac{i\tau[\gamma(\eta-1)-2]}{\gamma}$, $A_{4}=\frac{2i\tau(\gamma\eta-2)}{\gamma}$, and $A_{5}=\frac{2i\tau[\gamma(\eta-1)-1]}{\gamma}$. We consider the contour $(\eta,1)-(0,1/(1+\eta))-(0,0)-(1,0)-(1,1)-(\eta,1)$ around $U_{\eta}$. We compute
$TG|_{(\alpha,\gamma)=(\eta,1)}$ $=e^{i\tau(\eta-1)}+e^{i\tau(\eta+1)}-e^{i\tau(1-\eta)}=A,$
$B:=TG|_{(\alpha,\gamma)=(0,1/(1+\eta))}$ $=e^{-i\tau(1+\eta)}\frac{1-\eta}{1+\eta}+e^{i\tau(1+\eta)}\frac{2\eta}{1+\eta},$
$TG|_{(\alpha,\gamma)=(0,0)}$ $=e^{i\tau(\eta+1)}=C,$
$TG|_{(\alpha,\gamma)=(1,0)}$ $=e^{i\tau(\eta-1)}=D,$
$E:=TG|_{(\alpha,\gamma)=(1,1)}$ $=e^{i\tau(\eta-1)}\frac{2(1-\eta)}{2-\eta}+e^{-i\tau(\eta-3)}\frac{\eta}{2-\eta}.$
@@
$TG|_{\alpha=0}$ $=-e^{-i\tau(\eta+1-2/\gamma)}\frac{\gamma(\eta-1)}{\gamma(\eta+1)-2}+e^{i\tau(\eta+1)}$
$\quad+e^{i\tau(\eta+1-2/\gamma)}\frac{\gamma(\eta-1)}{\gamma(\eta+1)-2}$
(54) $=e^{i\tau(\eta+1)}+2(\eta-1)\frac{\sin[\tau(\eta+1-2/\gamma)]}{\eta+1-2/\gamma}\times i,$
@@
(55) $TG|_{\gamma=1}=e^{i\tau(\eta-1)}+\frac{2\eta}{\eta-2\alpha}\sin[\tau(\eta-2\alpha)]\times ie^{i\tau}.$
@@
(56) $TG|_{\alpha=1}=e^{i\tau(\eta-1)}+\frac{2\gamma\eta}{\gamma\eta-2}\sin[\tau(\eta-2/\gamma)]\times ie^{i\tau},$
@@ A.6.3. Surjectivity of $\phi_{3}$
$\phi_{3}^{-1}:P$ $\mapsto e^{i\tau(\eta+1)}-e^{i\tau(1-\eta)}+e^{i\tau(\eta-1)}=A,$
$Q$ $\mapsto e^{i\tau(\eta+1)}-1+1=e^{i\tau(\eta+1)}=C,$
(60) $R$ $\mapsto e^{i\tau(\eta+1)/3},$